d_{2}-METHOXY RADICAL DISPERSED FLUORESCENCE AND SEP SPECTROSCOPY.

\underline{XIAOYONG LIU}, VADIM STAKHURSKY, ERIC D. OLMON, VLADIMIR A. LOZOVSKY, TERRY A. MILLER, C.BRADLEY MOORE, The Ohio State University, Dept. of Chemistry, 120 W. 18th Avenue, Columbus, Ohio 43210, USA.

In the rotationally resolved LIF Ã ← X̃ spectrum of the jet-cooled d_{2}-methoxy radical (CHD_{2}O), combination bands of C-O (ν_{3}) and C-H (ν_{1}) stretches 1^{1}3^{1} and 1^{1}3^{2} were assigned. Analysis of vibrationally resolved Dispersed Fluorescence (DF) and rotationally resolved Stimulated Emission Pumping (SEP) spectra recorded upon excitation of these bands reinforces LIF spectral assignments and gives C-H and C-O stretch along with pseudo-Jahn-Teller active d_{2}-methyl rock (ν′′_{6}, ν′_{6}) vibrational frequencies in the X̃ state. A splitting of about ∼ 200 cm^{−1} for 1_{1}, 3_{1}1_{1}, 3_{2}1_{1} vibronic levels of CHD_{2}O is observed and may be due to the combination of vibronic degeneracy lifting and unquenched spin-orbit interaction. The DF spectra of CH_{3}O were recorded for excitation of Ã^{2}A_{1} 3^{n}4^{1} (n=1,2,3) combination bands containing asymmetric (ν_{4}) stretch. X̃^{2}E_{1} 3_{6}4_{1} and X̃^{2}E_{1} 3_{7}4_{1} rovibronic levels which lie ∼ 500 cm^{−1} and 1500 cm^{−1}, respectively, above the X̃^{2}E CH_{3}O barrier for dissociation CH_{2}O + H products, exhibit splitting of ∼ 120 cm^{−1}.